# arXiv:2603.02657v1[cs.RO] 03 Mar 2026

### v1

@@ IV-B1 Quantitative Analyses
TABLE I: Performance comparison under different obstacle densities.
(m) (%) (%) (m) (%) (%) (m) (%) (%) (m) (%) (%)
Ours 9.67 97.00 2.37 9.51 92.60 8.78 9.29 87.40 13.45 8.94 82.20 24.33
Ours w/o Virtual Obstacle Stage 6.19 51.60 13.18 5.88 45.80 20.88 5.42 39.40 29.67 5.06 30.08 36.15
@@
Traversal Capability and Survival: As expected, the Blind policy finds it almost impossible to navigate entire obstructed terrains, resulting in the lowest average distance to failure ($D$) and the lowest success rate ($S$). In contrast, Ours (full) maintains a near-perfect $D$ of approximately 9 m, demonstrating robust locomotion and sustained survival even when faced with heavily cluttered obstacles. Obstacle Avoidance Performance: The effectiveness of our two-stage RL approach is most evident in the step collision rate ($C$). Ours (full) achieves the lowest $C$ among all methods, effectively reducing unintended foot-obstacle collisions by 74% to 95% (depending on obstacle density) relative to the blind baseline. Conversely, the Blind policy possesses almost no capability for obstacle avoidance; as environmental density increases, almost every footstep inevitably results in a collision with the terrain debris.
@@ IV-B2 Detailed Ablation Analysis
We perform a detailed ablation analysis focusing on the two-stage RL curriculum, the reward formulation, and the exteroceptive modalities, as detailed in the lower section of Table I.
@@ IV-C Sim-to-Real Experiments
Sensor Noise and Map Modality Comparison: These low-profile objects represent the exact edge cases where traditional geometric proxies fail. In our real-world visualization, the physical depth sensor’s inherent noise floor completely obscures the millimeter-level height variations of these items, rendering the elevation map virtually indistinguishable from the flat ground. In stark contrast, the semantic map distinctly highlights these objects as high-cost hazards. This visual validation confirms the absolute necessity of semantic-geometry disentanglement for reliable real-world perception.
Semantic Avoidance: The core intelligence of SemLoco lies not merely in elevating the swing leg, but in its proactive foothold judgment. During experiment, if our policy predicts a foothold landing directly on a semantically hazardous object (e.g., a smartphone), the policy dynamically intervenes. It actively shortens the stride to secure a safe foothold immediately before the object. Subsequently, on the next gait cycle, it commands a high-clearance swing trajectory to cleanly step over the hazard. While a traditional blind controller might occasionally step over an object by pure luck if its foot happens to land perfectly in front of it, it fundamentally lacks this spatial judgment and proactive stride modulation.
Baseline Comparison: We qualitatively compare SemLoco against a blind baseline. The baseline policy exhibits a severe ”bulldozer” behavior: entirely oblivious to the surrounding fragility, it marches forward rigidly. Consequently, it collides with, kicks away, and aggressively tramples the scattered debris, posing a critical danger to both the robot and the objects. Conversely, SemLoco demonstrates significantly more deliberate navigation, effectively mitigating the majority of severe collisions with fragile debris. This highlights the potential of semantic-aware locomotion to improve operational safety in unstructured, real-world environments.
@@ V Conclusion
In this work, we introduced SemLoco, a semantic-aware locomotion framework that achieves precise foothold selection in dense, unstructured environments. By explicitly disentangling environmental semantics from geometric heights and employing a two-stage reinforcement learning curriculum, SemLoco successfully overcomes the geometric ambiguities and perceptual failure modes inherent to traditional elevation proxies. Extensive evaluations across both simulation and real-world indoor scenarios confirm that our approach significantly mitigates step collision rates, ensuring safe and deliberate foot placements without compromising overall locomotion efficiency.
Limitation and Future Work: While the proposed SemLoco framework significantly enhances semantic-aware navigation, simulation and deployment reveal limitations in both rigid-body dynamics and semantic representation. First, in terms of dynamics and hardware, extreme asymmetric footholds induce uncompensated angular momentum, occasionally resulting in observable yaw drift or pushing the joints near kinematic singularities. Additionally, aggressive clearance trajectories often exceed the physical bandwidth of the actuators, occasionally causing toe-stubbing due to tracking lag. Second, a noticeable sim-to-real gap stems from the semantic simulation environment. To manage computational overhead during RL training, complex real-world debris is approximated using simplified geometric primitives (e.g., rectangular grids). This approximation limits the policy’s exposure to realistic, high-fidelity object boundaries. Furthermore, our current pipeline compresses environmental semantics into a singular ”fragility” cost. Real-world applications, however, encompass open-vocabulary semantics that dictate diverse interactive behaviors. Future work will explore integrating high-fidelity semantic simulators and Vision-Language-Action (VLA) models to achieve comprehensive open-vocabulary semantic comprehension and generalized locomotion.

### v2

@@ IV-B1 Quantitative Analyses
Figure 5: Qualitative comparison of obstacle avoidance in a real cluttered environment.We compare our semantic-aware locomotion policy (Bottom) against an elevation-only baseline (Top). Elevation maps completely fail to capture small, low-profile objects (Fig 1). Consequently, Ours w/o Semantic Map degrades into a nearly blind policy, failing to adjust foot placements to avoid obstacles. As highlighted by the colored circles, the robot frequently steps on or kicks the scattered objects; the resulting environmental disruption is highly evident when comparing the Pre-Traversal and Post-Traversal states (the first and last columns). In contrast, our full framework leverages explicit semantic cues to achieve precise obstacle avoidance. The environment remains almost entirely undisturbed, with the exception of a negligible displacement of a power bank caused by an incidental grazing contact during the swing leg’s descent.
@@
Traversal Capability and Survival: As expected, the Blind policy finds it almost impossible to navigate entire obstructed terrains, resulting in the lowest average distance to failure ($D$) and the lowest success rate ($S$). In contrast, Ours (full) maintains a near-perfect $D$ of approximately 9 m, demonstrating robust locomotion and sustained survival even when faced with heavily cluttered obstacles.
Obstacle Avoidance Performance: The effectiveness of our two-stage RL approach is most evident in the step collision rate ($C$). Ours (full) achieves the lowest $C$ among all methods, effectively reducing unintended foot-obstacle collisions by 74% to 95% (depending on obstacle density) relative to the blind baseline. Conversely, the Blind policy possesses almost no capability for obstacle avoidance; as environmental density increases, almost every footstep inevitably results in a collision with the terrain debris.
@@ IV-B2 Detailed Ablation Analysis
We perform a detailed ablation analysis focusing on the two-stage RL curriculum, the reward formulation, and the exteroceptive modalities, as detailed in the Table I.
@@ IV-C Sim-to-Real Experiments
Sensor Noise and Map Modality Comparison: These small, low-profile objects represent the exact edge cases where traditional geometric proxies fail. In our visualization, the physical depth sensor’s inherent noise floor completely obscures the millimeter-level height variations of these items, rendering the elevation map virtually indistinguishable from the flat ground. In stark contrast, the semantic map distinctly highlights these objects as high-cost hazards. This visual validation confirms the absolute necessity of semantic-geometry disentanglement for reliable real-world perception.
Semantic Avoidance: The core intelligence of SemLoco lies not merely in elevating the swing leg, but in its proactive foothold judgment. During experiment, if our policy predicts a foothold landing directly on a semantically hazardous object (e.g., a smartphone), the policy dynamically intervenes. It actively shortens the stride or step off to the side to secure a safe foothold immediately before the object. Subsequently, on the next gait cycle, it commands a high-clearance swing trajectory to cleanly step over the hazard. While a traditional blind controller might occasionally step over an object by pure luck if its foot happens to land perfectly in front of it, it fundamentally lacks this spatial judgment and proactive stride modulation.
Baseline Comparison: We qualitatively compare SemLoco against a Ours w/o Semantic Map baseline. The baseline policy exhibits a severe ”bulldozer” behavior: entirely oblivious to the surrounding fragility, it marches forward rigidly. Consequently, it collides with, kicks away, and aggressively tramples the scattered debris, posing a critical danger to both the robot and the objects. Conversely, SemLoco demonstrates significantly more deliberate navigation, effectively mitigating the majority of severe collisions with fragile debris. This highlights the potential of semantic-aware locomotion to improve operational safety in unstructured, real-world environments.
@@ V Conclusion
In this work, we introduced SemLoco, a semantic-aware locomotion framework that achieves precise foothold selection in dense, unstructured environments. By explicitly disentangling environmental semantics from geometric heights and employing a two-stage reinforcement learning curriculum, SemLoco successfully overcomes the geometric ambiguities and perceptual failure modes inherent to traditional elevation proxies. Extensive evaluations across both simulation and real-world indoor scenarios confirm that our approach significantly mitigates step collision rates.
Limitation and Future Work: While the proposed SemLoco framework significantly enhances semantic-aware locomotion, simulation and deployment reveal limitations in both rigid-body dynamics and semantic representation. In terms of dynamics and hardware, extreme asymmetric footholds induce uncompensated angular momentum, occasionally resulting in observable yaw drift or pushing the joints near kinematic singularities. Additionally, it is extremely difficult for the real robots to replicate the foot trajectory and footswing height as seen in simulations. A noticeable sim-to-real gap stems from the semantic simulation environment. To manage computational overhead during training, complex real-world objects are approximated using simplified geometric primitives (e.g., cubes). This approximation limits the policy’s exposure to realistic object boundaries. Furthermore, our current pipeline compresses environmental semantics into a ”fragility” cost. However, real-world applications encompass open-vocabulary semantics that dictate diverse interactive behaviors. Future work will explore integrating high-fidelity semantic simulators and Vision-Language-Action (VLA) models to achieve comprehensive open-vocabulary semantic comprehension and generalized locomotion.